# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v2

@@
Fast Gradient Computation for RoPE Attention in Almost Linear Time
Yifang Chen yifangc@uchicago.edu. The University of Chicago.    Jiayan Huo jiayanh@arizona.edu. University of Arizona.    Xiaoyu Li 7.xiaoyu.li@gmail.com. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
The Rotary Position Embedding (RoPE) mechanism has become a powerful enhancement to the Transformer architecture, which enables models to capture token relationships when encoding positional information. However, the RoPE mechanisms make the computations of attention mechanisms more complicated, which makes efficient algorithms challenging. Earlier research introduced almost linear time, i.e., $n^{1+o(1)}$ where $n$ is the number of input tokens, algorithms for the forward computation under specific parameter settings. However, achieving a subquadratic time algorithm for other parameter regimes remains impossible unless the widely accepted Strong Exponential Time Hypothesis (SETH) is disproven. In this work, we develop the first almost linear time algorithm for backward computations in the RoPE-based attention under bounded entries. Our approach builds on recent advancements in fast RoPE attention computations, utilizing a novel combination of the polynomial method and the Fast Fourier Transform. Furthermore, we show that with lower bounds derived from the SETH, the bounded entry condition is necessary for subquadratic performance.
@@ 1 Introduction
The GPT-o3 [57], Llama 3.3 [54, 3], Claude 3.5 [4] are transformed-based Large Language Models (LLMs), have become important tools in natural language processing, which enables applications from machine translation to sentiment analysis. In the Transformer architecture, attention mechanisms, computationally intensive operations, compute token correlations within the sequence [72]. The efficiency of attention computations, both in forward computations and backward gradient computations, directly influenced the scalability and feasibility of training LLMs, especially when the size and input context length of these LLMs continue to grow [9, 6]. In recent research, Rotating Position Embedding (RoPE) [58] has become a popular modification to the attention mechanism, and it has enabled models to capture positional relationships between tokens with better expressiveness. The RoPE mechanism has been adopted in state-of-the-art models, such as Llama [70, 71, 54], Claude [4], Apple’s LLMs [29, 55], and many others, but the implementation of RoPE complicates attention computation due to the additional structure imposed by position-dependent rotations [58]. In recent work, [7] has demonstrated an efficient algorithm for forward computation of RoPE attention in the bounded entry regime. However, backward computation, the process of calculating gradients for model optimization, has been explored less.
@@
In Section 2, we present some relevant papers. In Section 3, we show essential components for the RoPE attention. In Section 4, we present the details of getting the gradients of RoPE-based attention mechanisms. Section 5 discusses the exact time complexity for computing RoPE attention gradients. Section 6 shows the fast computation of gradient using the polynomial method. Section 7 details the lower bounds of hardness. Finally, Section 8 provides conclusions and avenues for future work.
@@ 2 Related Work
In this paper, we study a variant of attention called RoPE attention. At a high level, RoPE gives more expressive power to the model in exchange for making the computational problem more complicated. In particular, many prior algorithms, such as the algorithm of [6], no longer apply to RoPE for fundamental reasons we will discuss. RoPE was proposed by [58] and has been used extensively in large-scale industrial models. Examples which are known to use RoPE include the open-source models released by Meta such as Llama [70] (see page 3), Llama 2 [71] (see page 5), Llama 3 [54] (see page 7), and the close-source LLM Claude 3.5 [4] released by Anthropic. Apple also incorporates RoPE into their LLM architecture (see [55], and page 3 of  [29]). The idea behind RoPE is to rotate the query and key vectors in the self-attention mechanism. The rotation is position-dependent and designed such that the inner product between two position-encoded vectors reflects their relative position in the input. In the RoPE attention mechanism, pairs of tokens with a longer relative distance will have a smaller correlation.
@@
The attention mechanism has often been criticized for its quadratic computational complexity concerning context length, a challenge that becomes more pronounced as the sequence length grows in today’s LLMs [2, 57, 54, 3, 5, 4]. However, this issue can be addressed using polynomial kernel approximation methods [1], which facilitate constructing the approximated attention matrix using low-rank approximations. Such techniques enable substantial improvements in computation speed, allowing a single attention layer to perform both training and inference nearly as fast as linear time [6, 8]. Our research further extends this efficiency to support multi-layer transformer architectures for both training and inference. In addition, these techniques can generalize to advanced attention mechanisms, such as tensor attention, while preserving the almost linear time complexity in both training and evaluation phases [9, 51]. Beyond this, alternative theoretical methods also exist. For example, the conv-basis approach introduced in [45] offers another avenue for speeding up attention computation.
@@
Theoretical Foundation of LLMs.
The attention has become a cornerstone in AI, particularly in large language models (LLMs), which excel in NLP tasks such as machine translation, text generation, and sentiment analysis due to their ability to capture complex contextual relationships. However, understanding the attention mechanism from a theoretical perspective remains an ongoing challenge. Several works have explored the theoretical foundations and computational complexities of attention [69, 80, 17, 6, 66, 18, 20, 43, 56, 68, 12, 22, 16, 65, 10], focusing on areas such as efficient attention [31, 59, 67, 40, 45, 51, 49, 47, 46, 21, 44, 35, 34, 30, 73, 32, 74, 36, 8, 24, 59, 41, 50, 60, 61], optimization [23, 52, 62, 27, 53, 26, 25], and the analysis of emergent abilities [15, 77, 11, 37, 79, 42, 78, 19, 48, 33, 76, 63, 75, 78, 38]. Notably, [80] introduced an algorithm with provable guarantees for attention approximation, [39] proved a lower bound for attention computation based on the Strong Exponential Time Hypothesis, and [6] provided both an algorithm and hardness results for static attention computation using low-rank approximation [28, 64].
@@ 3 Preliminaries on RoPE Attention
In Section 3.1, we talk about the notation and foundational concepts. In Section 3.2, we formalize our problems. In Section 3.3, we talk about polynomial approximation of the exponential function. In Section 3.4, we talk about the time complexity of matrix multiplications, setting up the framework for analyzing efficiency in attention computation. In Section 3.5, we talk about the Strong Exponential Time Hypothesis (SETH). In Section 3.6, we talk about mathematical properties and tricks, such as the tensor trick and row-wise Kronecker products, which enable efficient matrix-vector operations. In Section 3.7, we give definitions of our key terms. In Section 3.8, we talk about the reformulation of the loss function using the tensor trick and analyze the computational complexity of the reformulated expressions.
@@ 3.2 Problem Definition
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. Given three $n\times d$ matrices $Q,K,V$ with the guarantee that $\|Q\|_{\infty},\|K\|_{\infty},\|V\|_{\infty}\leq B$ and $\|W\|_{\infty}\leq 1$. We define matrix $A\in\mathbb{R}^{n\times n}$ as, for $i,j\in[n]$,
$A_{i,j}:=\exp(Q_{i,*}W_{i-j}K_{j,*}^{\top}/d).$
We define $D:=\operatorname{diag}(A{\bf 1}_{n})$. The goal of General Approximate RoPE Attention Computation is to output a matrix $T\in\mathbb{R}^{n\times d}$ such that $\|T-\mathsf{ARAttC}\|_{\infty}\leq\epsilon$ is small, where $\mathsf{ARAttC}:=D^{-1}AV$. For matrix $M$, we use $\|M\|_{\infty}:=\max_{i,j}|M_{i,j}|$. Note that the $1/d$ factor inside $\exp$ in the definition of $A$ is a normalization factor.
@@
$A(X_{1},X_{2})_{i,j}:=$ $~{}\exp((A_{1}X_{1})_{i,*}W_{i-j}(A_{2}X_{2})^{\top}_{j,*}/d)=~{}\exp(A_{1,i,*}X_{1}W_{i-j}X_{2}^{\top}A^{\top}_{2,j,*}/d).$
We define $w_{i-j}:=\operatorname{vec}(W_{i-j})\in\mathbb{R}^{d^{2}}$ and define $\mathsf{W}$ such that $\mathsf{W}_{j_{0},*}$ is an $1\times d^{2}$ block and $\mathsf{W}_{i+(j-1)n}:=w_{i-j}^{\top}$. Here, let $\mathsf{A}:=A_{1}\otimes A_{2}\in\mathbb{R}^{n^{2}\times d^{2}}$ and $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We can show that
$A_{1,i,*}X_{1}W_{i-j}X_{2}^{\top}\mathsf{A}^{\top}_{2,j,*}=$ $~{}(A_{1,i,*}\otimes A_{2,j,*})(X_{1}\otimes X_{2})\operatorname{vec}(W_{i-j})=~{}\mathsf{A}_{i+(j-1)n,*}\mathsf{X}w_{i-j},$
@@
$A(\mathsf{X})_{i,j}=$ $~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes w_{i-j}^{\top})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})=~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes\mathsf{W}_{i+(j-1)n,*})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1}).$
@@
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. For $i,j\in[n]$, let $\mathsf{W}\in\mathbb{R}^{n^{2}\times d^{2}}$ such that $\mathsf{W}_{i+(j-1)n,*}=\operatorname{vec}({W_{i-j}})$. Here, we suppose four $n\times d$ matrices $A_{1},A_{2},A_{3},E$, and we have three $d\times d$ matrices $X_{1},X_{2},Y$. Let $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We define the matrix $A(\mathsf{X})\in\mathbb{R}^{n\times n}$ as the matrix representation of $\exp(\underbrace{(\mathsf{A}\oslash\mathsf{W})}_{n^{2}\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})$ and the $n\times n$ matrix $D(\mathsf{X}):=\operatorname{diag}(\underbrace{A(\mathsf{X})}_{n\times n}\underbrace{{\bf 1}_{n}}_{n\times 1}).$ The RoPE attention optimization problem is formulated as follows:
$\min_{\mathsf{X}\in\mathbb{R}^{d^{2}\times d^{2}}}\mathsf{Loss}(\mathsf{X}):=\min_{\mathsf{X}\in\mathbb{R}^{d^{2}\times d^{2}}}0.5\|D(\mathsf{X})^{-1}A(\mathsf{X})A_{3}Y-E\|_{F}^{2}.$
@@
$\frac{\mathrm{d}\mathsf{Loss}(X_{1},X_{2})}{\mathrm{d}X_{1}}=$ $~{}\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}\frac{\mathrm{d}\mathsf{X}}{\mathrm{d}X_{1}}=~{}\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}\frac{\mathrm{d}(X_{1}\otimes X_{2})}{\mathrm{d}X_{1}}=~{}\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}(I_{d\times d}\otimes X_{2}).$

### v3

# RoPE Attention Can Be Trained in Almost Linear Time

@@
RoPE Attention Can Be Trained in Almost Linear Time
Yang Cao and Jiayan Huo and Yingyu Liang and Zhenmei Shi and Zhao Song Wyoming Seminary.University of Arizona.The University of Hong Kong.University of Wisconsin-Madison.magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.
The Rotary Position Embedding (RoPE) mechanism has become a powerful enhancement to the Transformer architecture, which enables models to capture token relationships when encoding positional information. However, the RoPE mechanisms make the computations of attention mechanisms more complicated, which makes efficient algorithms challenging. Earlier research introduced almost linear time algorithms for the forward computation under specific parameter settings of bounded entries (i.e., in time $n^{1+o(1)}$ where $n$ is the number of input tokens), but has not addressed backward computation. In this work, we develop the first almost linear time algorithm for backward computations in the RoPE-based attention under bounded entries. Our approach builds on recent advancements in fast RoPE attention computations, utilizing a novel combination of the polynomial method and the Fast Fourier Transform. Furthermore, we show that with lower bounds derived from the Strong Exponential Time Hypothesis (SETH), the bounded entry condition is necessary for subquadratic performance.
@@ 1 Introduction
The GPT-o3 [22], Llama 3.3 [20, 3], Claude 3.5 [9] are transformed-based Large Language Models (LLMs), have become important tools in natural language processing, which enables applications from machine translation to sentiment analysis. In the Transformer architecture, attention mechanisms, computationally intensive operations, compute token correlations within the sequence [26]. The efficiency of attention computations, both in forward computations and backward gradient computations, directly influenced the scalability and feasibility of training LLMs, especially when the size and input context length of these LLMs continue to grow [6, 4]. In recent research, Rotating Position Embedding (RoPE) [23] has become a popular modification to the attention mechanism, and it has enabled models to capture positional relationships between tokens with better expressiveness. The RoPE mechanism has been adopted in state-of-the-art models, such as Llama [24, 25, 20], Claude [9], Apple’s LLMs [14, 21], and many others, but the implementation of RoPE complicates attention computation due to the additional structure imposed by position-dependent rotations [23]. In recent work, [5] demonstrated an efficient algorithm for forward computation of RoPE attention in the bounded entry regime. However, backward computation, the process of calculating gradients for model optimization, has been explored less.
@@
In Section 2, we present some relevant papers. In Section 3, we show essential components for the RoPE attention. Section 4 gives the closed form of the RoPE Attention gradient and discusses the exact time complexity to compute it. Section 5 shows the fast computation of the RoPE Attention gradient in almost linear time. Section 6 details the lower bounds of hardness. Finally, Section 7 provides conclusions and avenues for future work.
@@ 2 Related Work
At a high level, RoPE gives more expressive power to the model in exchange for making the computational problem more complicated. In particular, many prior algorithms, such as the algorithm of [4], no longer apply to RoPE for fundamental reasons we will discuss. RoPE was proposed by [23] and has been used extensively in large-scale industrial models. Examples which are known to use RoPE include the open-source models released by Meta such as Llama [24] (see page 3), Llama 2 [25] (see page 5), Llama 3 [20] (see page 7), and the close-source LLM Claude 3.5 [9] released by Anthropic. Apple also incorporates RoPE into their LLM architecture (see [21], and page 3 of  [14]).
@@
The attention mechanism has often been criticized for its quadratic computational complexity concerning context length, a challenge that becomes more pronounced as the sequence length grows in today’s LLMs [1, 22, 20, 3, 8, 9]. However, this issue can be addressed using polynomial kernel approximation methods [2], which facilitate constructing the approximated attention matrix using low-rank approximations. Such techniques enable substantial improvements in computation speed, allowing a single attention layer to perform both training and inference nearly as fast as linear time [4, 7]. [18] further extends this efficiency to support multi-layer transformer architectures for both training and inference. In addition, these techniques can generalize to advanced attention mechanisms, such as tensor attention, while preserving the almost linear time complexity in both training and evaluation phases [6, 19]. Beyond this, alternative theoretical methods also exist. For example, the conv-basis approach introduced in [17] offers another avenue for speeding up attention computation.
@@ 3 Preliminaries on RoPE Attention
In Section 3.1, we talk about the notation and foundational concepts. In Section 3.2, we formalize our problems. In Section 3.3, we talk about the reformulation of the loss function using the tensor trick and analyze the computational complexity of the reformulated expressions.
@@ 3.2 Problem Definition
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. Given three $n\times d$ matrices $Q,K,V$ with the guarantee that $\|Q\|_{\infty},\|K\|_{\infty},\|V\|_{\infty}\leq B$ and $\|W\|_{\infty}\leq 1$. We define matrix $A\in\mathbb{R}^{n\times n}$ as, for $i,j\in[n]$, $A_{i,j}:=\exp(Q_{i,*}W_{i-j}K_{j,*}^{\top}/d).$ We define $D:=\operatorname{diag}(A{\bf 1}_{n})$. The goal of General Approximate RoPE Attention Computation is to output a matrix $T\in\mathbb{R}^{n\times d}$ such that $\|T-\mathsf{ARAttC}\|_{\infty}\leq\epsilon$ is small, where $\mathsf{ARAttC}:=D^{-1}AV$. For matrix $M$, we use $\|M\|_{\infty}:=\max_{i,j}|M_{i,j}|$. Note that the $1/d$ factor inside $\exp$ in the definition of $A$ is a normalization factor.
@@
$A(X_{1},X_{2})_{i,j}:=$ $~\exp((A_{1}X_{1})_{i,*}W_{i-j}(A_{2}X_{2})^{\top}_{j,*}/d)$
$=$ $~\exp(A_{1,i,*}X_{1}W_{i-j}X_{2}^{\top}A^{\top}_{2,j,*}/d).$
We define $w_{i-j}:=\operatorname{vec}(W_{i-j})\in\mathbb{R}^{d^{2}}$ and define $\mathsf{W}$ such that $\mathsf{W}_{j_{0},*}$ is an $1\times d^{2}$ block and $\mathsf{W}_{i+(j-1)n,*}:=w_{i-j}^{\top}$. Here, let $\mathsf{A}:=A_{1}\otimes A_{2}\in\mathbb{R}^{n^{2}\times d^{2}}$ and $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We can show that
$~A_{1,i,*}X_{1}W_{i-j}X_{2}^{\top}\mathsf{A}^{\top}_{2,j,*}$
$=$~ $\mathsf{A}_{i+(j-1)n,*}\mathsf{X}w_{i-j},$
@@
$A(\mathsf{X})_{i,j}=$ $~\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes w_{i-j}^{\top})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})$
$=$ $~\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes\mathsf{W}_{i+(j-1)n,*})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1}).$
@@
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. For $i,j\in[n]$, let $\mathsf{W}\in\mathbb{R}^{n^{2}\times d^{2}}$ such that $\mathsf{W}_{i+(j-1)n,*}=\operatorname{vec}({W_{i-j}})$. Here, we suppose four $n\times d$ matrices $A_{1},A_{2},A_{3},E$, and we have three $d\times d$ matrices $X_{1},X_{2},Y$. Let $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We define the matrix $A(\mathsf{X})\in\mathbb{R}^{n\times n}$ as the matrix representation of $\exp(\underbrace{(\mathsf{A}\oslash\mathsf{W})}_{n^{2}\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})$ and the $n\times n$ matrix $D(\mathsf{X}):=\operatorname{diag}(\underbrace{A(\mathsf{X})}_{n\times n}\underbrace{{\bf 1}_{n}}_{n\times 1}).$ The RoPE attention optimization problem $\min_{\mathsf{X}\in\mathbb{R}^{d^{2}\times d^{2}}}\mathsf{Loss}(\mathsf{X})$ is formulated as follows:
$\min_{\mathsf{X}\in\mathbb{R}^{d^{2}\times d^{2}}}0.5\|D(\mathsf{X})^{-1}A(\mathsf{X})A_{3}Y-E\|_{F}^{2}.$
@@
$\frac{\mathrm{d}\mathsf{Loss}(X_{1},X_{2})}{\mathrm{d}X_{1}}=$ $~\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}\frac{\mathrm{d}\mathsf{X}}{\mathrm{d}X_{1}}$
$=$ $~\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}\frac{\mathrm{d}(X_{1}\otimes X_{2})}{\mathrm{d}X_{1}}$
$=$ $~\frac{\mathrm{d}\mathsf{Loss}(\mathsf{X})}{\mathrm{d}\mathsf{X}}(I_{d\times d}\otimes X_{2}).$